OpenStudy (anonymous):

Heather's family bought tickets at a local carnival. Her brother took half the number of tickets that her cousin took, which was one-third the number Heather took. If her brother took ten tickets, how many tickets did Heather take? twenty thirty forty sixty

OpenStudy (anonymous):

heather bought say \(n\) tickets, the her cousin bought \(\frac{1}{3}n\) tickets

OpenStudy (anonymous):

Thats clear to me.

OpenStudy (anonymous):

her brother bought have of the amount her cousin bought i.e. \[\frac{1}{2}\times \frac{1}{3}n=\frac{1}{6}n\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

since her brother bought \(10\) tickets you know \[\frac{1}{6}n=10\]

OpenStudy (anonymous):

yes so I divided that on a calculator which is 1.666666667

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

so \(n=6\times 10=60\)

OpenStudy (anonymous):

ok well im good

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

if one sixth of a number is ten, that number must be sixty

OpenStudy (anonymous):

@satellite73 yes so now..

OpenStudy (anonymous):

@satellite73 thank you so much